OpenStudy (anonymous):

solve formula for variable k 3/7+3y=k/u

OpenStudy (anonymous):

\[u*(\frac{3}{7}+3y)=k\] \[\frac{3u}{7}+3uy=k\]

OpenStudy (anonymous):

\[3/7+3y=k/u\] \[u(3/7+3y)=k\] \[u((3+21y)/7)=k\] \[3u+63y/7=k\]